# Self-Quenching Effect of the Decay of Localized Surface Plasmons: Classical and Quantum Perspectives

### v1

@@ 1 Introduction
Nanometallic (plasmonic) structures exhibit a remarkable ability to confine and concentrate electromagnetic (EM) field energy within subwavelength volumes, making them essential components in modern nanophotonics [barnes2003surface, schuller2010plasmonics, novotny2012principles, alekseeva2019single, nam2019chemical]. Their ability to create highly localized and intense optical fields has prompted a fundamental reevaluation of light–matter interactions at the nanoscale. The field of plasmonics offers a rich platform for manipulating light at nanometer scales, with applications spanning nanophotonic lasers [gaio2019nanophotonic], optical amplifiers [qin2024quantum], metamaterials [cai2010optical], biochemical sensing [aslan2005plasmon, jain2007review, shalabney2011sensitivity], and nanoantennas [monticone2017optical].
@@
Recent research has increasingly focused on the quantum properties of localized surface plasmons (LSPs) [jacob2011plasmonics, van2012spontaneous, scholl2012quantum, bozhevolnyi2017plasmonics, kolwas2019decay]. In conventional cavity quantum electrodynamic (CQED) studies, the spontaneous emission of a quantum emitter is enhanced by its placement within an external optical cavity whose spectral and spatial characteristics modify the vacuum fluctuations [haroche1989cavity, walther2006cavity, guzatov2012plasmonic, hugall2018plasmonic, franke2019quantization].
@@
In plasmonic CQED (PCQED), principles of cavity quantum electrodynamics are implemented to nanoscale plasmonic cavities, in which surface plasmons confine electromagnetic fields to subwavelength volumes with strongly enhanced field intensities. Within these nanostructured electromagnetic environments, the interaction dynamics with an atom- or molecule-like emitter is determined not only by the cavity quality factor, but also by the extremely small effective mode volume [kristensen2014modes, ginzburg2016cavity, hugall2018plasmonic, wei2021plasmon]. Such configurations enable the redistribution of the local density of states (LDOS), leading to either the enhancement or suppression of spontaneous emission on ultrafast time scales via the Purcell effect.
This study explores a fundamentally different regime: the transient free decay of LSPs’ excitations in a single metal nanoparticles (NPs). The LSP resonant excitations are modeled as plasmonic quasi-particles (PQPs) that function as the transient emitters of the TM (transverse magnetic) fields, forming a self-created resonant near-field nano-cavity of confined radial fields coupled to the oscillating surface charges. The radially polarized electric fields hold particular significance for light–matter interactions in metal NPs, in the spectral ranges where the real component of the dielectric function is negative. We consider an alteration of a decay process by a structured near-field environment, resulting in the modification of the LDOS available for the plasmonic emitter, which introduces the self-quenching effect due to the intrinsic nonradiative losses in a metal NP.
The quantum modeling developed here is grounded in classical electrodynamics, which provides the eigenfrequencies and damping rates of quasi-normal modes (QNM) of LSPs via dispersion relations for surface-bound TM modes in spherical geometries [fuchs1992basic, kolwas1996optical, kolwas2006smallest]. This classical foundation enabled the construction of a discrete-level quantum many-electron system versus NPs size, allowing for a consistent interpretation of the decay of LSPs excitations in the quasi-particles framework. Overall, the study offers a unified view that bridges classical and quantum perspectives and yields insights into the decay dynamics specific to LSP excitations.
This paper is organized as follows: Sections 2 and 3 revisit crucial aspects of both classical and quantum foundations of localized surface plasmon (LSP) excitations in metal nanoparticles, necessary for further analysis. Section 4 introduces the fundamentals of spontaneous emission and describes how the Purcell effect results from modified electromagnetic environments. Section 5 presents the central concept of the dual emitter-cavity nature of PQP’s decay dynamics. It starts with a classical description of LSP resonances and their relation to a quantum model. It discusses factors influencing the quality factor, the density of optical states associated with confined fields of LSP excitations, and a self-consistent extension of Fermi’s Golden Rule for a plasmonic emitter. Finally, Section 6 presents concluding remarks, including a discussion of the physical implications and the specific advantages associated with the investigation of plasmonic nanostructures.
2 Maxwell Picture of Localised Surface Plasmons: Lorenz–Mie Scattering vs. Dispersion Relation
@@
In classical electrodynamics, the response of a spherical nanoparticle (NP) of radius $R$ to an external electromagnetic (EM) wave is described by solving Maxwell’s equations with boundary conditions at the NP surface. The EM fields inside and outside the NP are typically expanded using Lorenz–Mie scattering theory as an infinite series of spherical multipole modes with angular momenta $l=1,2,3,\ldots$ [born1999principles]. The general solution is a sum of transverse magnetic (TM) and transverse electric (TE) modes.
@@ 2.1 Dispersion Relation for TM Modes
The dispersion relation for QNMs associated with LSPs is derived from the source-free Mie formalism [mie1908], as presented in classical works [stratton2007electromagnetic, born1999principles, ruppin1982electromagnetic, fuchs1992basic]. Scalar solutions to the Helmholtz equation yield two distinct vectorial modes: TE modes: $E_{r}=0$, and TM modes: $E_{r}\neq 0$, which alone couple to surface charge oscillations.
The complex frequencies of QNMs are roots of the dispersion relation for TM modes with angular momentum $l$, derived from the continuity of EM field components at the interface $r=R$:
@@
Unlike surface plasmon polaritons (SPPs), for which DR yields complex wave vectors for real frequencies, spherical LSP modes require solving for complex eigenfrequencies:
@@
where $\omega_{l}(R)$ is the resonance frequency and $\Gamma_{l}(R)$ is the total damping rate of LSP oscillations.
@@
Figure 1a illustrates the dependence $\omega_{l}(R)$ and $\Gamma_{l}(R)$ for gold nanospheres for consecutive radii (data from [kolwas2013damping]), retardation included).
Figure 1: Classical and quantum representations of LSP dynamics. (a) Resonant frequencies $\hbar\omega_{l}(R)$ and damping rates $\hbar\Gamma_{l}(R)$ of QNMs as a function of radius $R$, including retardation effects [kolwas2013damping]. (b) Mapping multipolar modes $l$ to energy levels of a quantum plasmonic quasi-particle (PQP), modeled as independent two-level systems $S_{l}$.
Spatial field confinement stems from the negative real part of $\varepsilon_{in}(\omega)$ (anomalous dispersion) in a dielectric medium ($\varepsilon_{out}>0$), which causes the radial electric field to decay evanescently. Outside the NP, radial components fall off as $1/r^{2}$, and tangential components as $1/r$ [fuchs1992basic].
The dispersion relation approach has been extensively studied [derkachova2016dielectric, kolwas2017modification, kolwas2020impact] to determine $\omega_{l}(R)$ and $\Gamma_{l}(R)$ for a wide range of radii and materials.
2.2 Damping Rates from the Dispersion Relation
@@
where $\Gamma_{l}^{r}(R)$ is the radiative damping rate (from the collisionless model), and $\Gamma^{nr}(R)=\gamma(R)/2$, if assumed to account for electron collisions only, with:
@@
This effective collision rate appears explicitly in $\varepsilon_{in}$, incorporating both bulk ohmic losses $\gamma_{b}$ and the size-dependent contribution from surface scattering of electrons, where $v_{F}$ denotes the Fermi velocity and $C\approx 1$ for noble metals such as Au and Ag (e.g., [derkachova2016dielectric, kolwas2020impact]). The data from [derkachova2016dielectric] (Figure 2) also phenomenologically incorporate the influence of interband transitions on the polarizability. However, the corresponding transition rate is not an explicit parameter in the model of the dielectric function.
Figure 2: (a) Total damping rates $\Gamma_{l}(R)$ (for l=16l=1-6) from DR. (b) Close-up view showing that $\Gamma_{l}(R)\leq\gamma(R)/2$, suggesting non-additivity of radiative and nonradiative terms. Data from [kolwas2013damping].
@@
which may stem from a suppression of the nonradiative damping [sonnichsen2002drastic, kolwas2013damping]. In particular, Sonnichsen et al. [sonnichsen2002drastic] attributed the suppression of nonradiative damping primarily to a reduction in interband transition processes.
@@
3 Foundations for Describing LSPs in Metal Nanoparticles as Quantum Plasmonic Quasi-Particle
The concept of modeling localized surface plasmon (LSP) excitations as quasi-particles arises from their resonant behavior followed by relaxation, observable through the spectral properties of scattered or absorbed light. This approach establishes a bridge between the classical electromagnetic description based on Maxwell’s equations and a quantum-mechanical framework where plasmonic excitations are treated as quantized states of a collective many-electron system.
@@ 3 Foundations for Describing LSPs in Metal Nanoparticles as Quantum Plasmonic Quasi-Particle
Each multipolar mode $l$ corresponds to a discrete energy level of a quantum plasmonic quasi-particle (PQP), modeled as a set of uncoupled two-level subsystems $S_{l}$ undergoes transient decay to the ground state $|0\rangle$ after excitation.
@@
3.1 Closed System. Plasmonic Quasiparticles and Plasmonic Cavities Conceptualized as Equivalent Quantum Systems
In conventional CQED, models such as the Jaynes-Cummings Hamiltonian describe the interaction between a two-level atom and a single quantized mode of an external cavity field [jaynes1963comparison]. This model assumes a closed system that conserves the number of quanta.
In the specific case of a closed confined subsystem $S_{l}$ of POP, it forms an integrated quantum system in which quantized TM polarized fields sourced in LSP oscillations and quasi-particle dynamics are intrinsically coupled (see Figure 3). Each surface-cavity mode $l$ is described using an anti-normally ordered Hamiltonian:
@@
Such a self-consistent model provides a compact, physically grounded quantum framework for describing many-electron systems performing coherent collective motion. These systems can be viewed as micro- and macroscopic quantum structures (from tens to hundreds of nanometers in size), sharing characteristics related to the phenomenon of macroscopic quantum coherence demonstrated in superconducting circuits [devoret1985measurements, martinis1985energy] and recently awarded by the Nobel Prize.
3.2 Decay dynamics of plasmonic quasi-particles
In realistic conditions, the plasmonic quasi-particle system $S=\sum_{l}S_{l}$ interacts with its environment $E$, forming an open quantum system. Environmental coupling introduces dissipation and decoherence, leading to energy loss and phase randomization. These effects can be approximately described within the Lindblad master equation formalism under the Markov approximation [breuer2002theory].
@@ 3.2 Decay dynamics of plasmonic quasi-particles
The system evolves through two main decay channels: radiative decay, associated with photon emission, modeled by $L_{r,l}=\sqrt{2\Gamma_{l}^{reff}}\,\sigma_{l}^{-}$ and nonradiative decay, associated with internal losses (ohmic heating, electron–electron or electron–surface scattering), modeled by $L_{nr,l}=\sqrt{2\Gamma^{nr}}\,\sigma_{l}^{-}$. The radius dependence of the rates $\Gamma_{l}^{reff}(R)$ and $\Gamma^{nr}(R)$ follows from the classical modeling.
Solving the Lindblad equation for POP [kolwas2019decay, kolwas2023optimization] yields exponentially decaying populations and coherences:
@@
4 Concept of Spontaneous Emission
4.1 Fundamental Concepts of Spontaneous Emission
@@
where $\omega_{21}$ is the transition frequency and $d_{12}$ is the dipole matrix element. This expression is sometimes analogously applied to describe the radiative damping of dipolar LSP modes (e.g. [moustafa2023bandwidth]).
When the emitter is placed in an inhomogeneous medium, such as an optical cavity, the spontaneous emission rate is modified by the available EM states of the surrounding environment, altering the vacuum field mode structure and the local density of states. As a result, the emission rate can be enhanced or suppressed compared to that in free space (e.g., [martin2014quantum, novotny2012principles]. This fundamental sensitivity of spontaneous emission to the electromagnetic environment forms the basis of CQED (e.g., [maier2006effective, kristensen2012generalized, derom2012resonance, sauvan2013theory, pelton2015modified]) and underpins emission control in structured photonic media.
4.2 Cavity LDOS and the Purcell Effect
Cavity quantum electrodynamics (CQED) and its plasmonic counterpart (PQED) investigate how optical and plasmonic cavities affect spontaneous emission via the Purcell effect, which quantifies enhancement (or suppression) of emission due to increased LDOS near the emitter’s resonance [haroche1989cavity, walther2006cavity, hugall2018plasmonic].
@@ 4.2 Cavity LDOS and the Purcell Effect
For the conventional emitters in a plasmonic cavity, strong field confinement leads to very small $V_{eff}$, enabling high Purcell enhancement despite low $Q$. Plasmonic cavities of different geometries (e.g., bowtie, gap antenna) are essential tools in nanophotonics for squeezing light down to the nanoscale to unlock powerful light-matter interactions. However, in proximity to metallic surfaces, nonradiative decay processes become predominant, leading to quenching of the emission and a concomitant decrease in the lifetime of the emitter (e.g. [lakowicz2005radiative, anger2006enhancement, bharadwaj2006nanoplasmonic, guzatov2012plasmonic, olejnik2012plasmonic, rasskazov2018plasmon, kongsuwan2018suppressed, hugall2018plasmonic, hildebrandt2023plasmonic])
5 The Dual Functionality of Plasmonic Quasi-Particles

### v2

@@ 1 Introduction
Nanometallic (plasmonic) structures exhibit a remarkable ability to confine and concentrate electromagnetic (EM) field energy within subwavelength volumes, making them essential components of modern nanophotonics 4, 55, 46, 1, 45. Their ability to create highly localized and intense optical fields has prompted a fundamental reevaluation of light–matter interactions at the nanoscale. The field of plasmonics provides a rich platform for manipulating light at nanometer scales, with applications ranging from nanophotonic lasers 15 and optical amplifiers 49 to metamaterials 9, biochemical sensing 3, 23, 57, and optical nanoantennas 43.
@@
Recent research has increasingly focused on the quantum aspects of localized surface plasmons (LSPs) 22, 61, 54, 7, 31. In conventional cavity quantum electrodynamics (CQED), the spontaneous emission of a quantum emitter can be strongly modified by placing it inside an optical cavity whose spectral and spatial characteristics alter the vacuum fluctuations of the electromagnetic field 19, 62, 18, 21, 13.
@@
In plasmonic implementations of CQED (PCQED), nanoscale metallic cavities confine electromagnetic fields to extremely small effective volumes while maintaining significant field intensities. In such systems, the interaction dynamics between the cavity field and an atom- or molecule-like emitter is determined not only by the cavity quality factor but also by the strongly reduced effective mode volume 34, 17, 21, 63. The resulting modification of the local density of optical states (LDOS) typically leads to substantial enhancement or suppression of spontaneous emission rates via the Purcell effect.
However, conventional approaches typically treat the quantum emitter and the cavity as distinct, independent subsystems. The present work addresses a fundamentally different and highly original physical scenario: the decay dynamics of a plasmonic excitation that acts simultaneously as the emitter and as the source of its own structured electromagnetic environment. Here, the plasmon itself is described as a plasmonic quasi-particle (PQP) — a collective, bosonic many-electron excitation embedded within a self-created resonant near-field nanocavity of confined radial fields.
A cornerstone of standard decay models of atom-like emitters is the assumption of additivity — namely, that the total damping rate of an excitation can be expressed as the sum of independent radiative and nonradiative decay channels. In the case of localized surface plasmons (LSP) excitations, we show that this assumption does not generally hold. Within the self-generated electromagnetic environment of a plasmonic nanoparticle, intrinsic nonradiative (ohmic) losses modify the local density of optical states (LDOS) of the confined field. This dynamic coupling introduces a feedback mechanism that alters the effective radiative decay rate of the plasmonic excitation. As a result, the radiative and nonradiative decay channels become intrinsically coupled and cannot, in general, be treated as independent contributions.
To capture this effect, the theoretical framework developed here bridges classical Maxwell electrodynamics with a quantum description of plasmonic excitations. The classical dispersion relations for surface-bound transverse-magnetic modes, with surface collisions and retardation effects fully included, provide the eigenfrequencies and damping rates of the quasi-normal modes supported by spherical nanoparticles 14, 26, 27. These classical results serve as the foundation for constructing a discrete quantum description of LSP excitations and for analyzing their decay dynamics within a self-consistent emitter–cavity system.
On this basis, we derive a self-consistent extension of Fermi’s Golden Rule for plasmonic quasi-particles that incorporates the coupling between the emission process and the self-generated cavity impact. This formulation yields an analytical expression for the total damping rates, revealing a size-dependent reduction in the total damping rate, particularly prominent in larger size ranges and higher multipolarity modes. This intrinsic suppression of radiative emission—referred to here as the self-quenching effect — arises directly from the nonadditive coaction of radiative and nonradiative channels.
Ultimately, the resulting framework provides a unified perspective that connects classical electrodynamics to the quantum bosonic characteristics of plasmonic quasi-particles, offering novel, analytically tractable insights into the fundamentally coupled decay dynamics of localized surface plasmons.
2 Classical Maxwell Foundations
@@
In classical electrodynamics, the response of a spherical nanoparticle (NP) of radius $R$ to an external electromagnetic (EM) wave is described by solving Maxwell’s equations with boundary conditions at the NP surface. The EM fields inside and outside the NP are typically expanded using Lorenz–Mie scattering theory as an infinite series of spherical multipole modes with angular momenta $l=1,2,3,\ldots$ 6. The general solution is a sum of transverse magnetic (TM) and transverse electric (TE) modes.
@@ 2.1 Dispersion Relation for TM Modes
The dispersion relation for QNMs associated with LSPs is derived from the source-free Mie formalism 42, as presented in classical works 60, 6, 52, 14. Scalar solutions to the Helmholtz equation yield two distinct vectorial modes: TE modes: $E_{r}=0$, and TM modes: $E_{r}\neq 0$, which alone couple to surface charge oscillations.
The complex frequencies of QNMs are roots of the dispersion relation for TM modes with angular momentum $l$, derived from the continuity of EM field components at the interface $r=R$ 14, 26, 27:
@@
Unlike the commonly employed driven-propagation spectral framework for planar surface plasmon polaritons (SPPs), in which the dispersion relation yields complex wave vectors at real excitation frequencies, the quasi-normal mode description of LSPs on spherical metallic nanoparticles necessitates the allowance of complex eigenfrequencies. This requirement is analogous to the complex-frequency eigenmode formulation used for planar SPPs 38:
@@
where $\omega_{l}(R)$ is the resonance frequency and $\Gamma_{l}(R)$ is the total damping rate of LSP oscillations of QNM TM modes.
@@
Figure 1a illustrates the dependence $\omega_{l}(R)$ and $\Gamma_{l}(R)$ for gold nanospheres for consecutive radii (data from 28), retardation included).
Figure 1: Classical and quantum representations of LSP dynamics. (a) Resonant frequencies $\hbar\omega_{l}(R)$ and damping rates $\hbar\Gamma_{l}(R)$ of QNMs as a function of radius $R$, including retardation effects 28. (b) Mapping multipolar modes $l$ to energy levels of a quantum plasmonic quasi-particle (PQP), modeled as independent two-level systems $S_{l}$.
Spatial field confinement stems from the negative real part of $\varepsilon_{in}(\omega)$ (anomalous dispersion) in a dielectric medium ($\varepsilon_{out}>0$), which causes the radial electric field to decay evanescently. Outside the NP, radial components fall off as $1/r^{2}$, and tangential components as $1/r$ 14.
The dispersion-relation approach has been the subject of our extensive studies employing a size-dependent dielectric function $\varepsilon_{in}(\omega,R)$, which explicitly accounts for surface-collision effects arising from finite nanoparticle sizes as well as retardation effects in gold and silver nanoparticles embedded in various dielectric environments 10, 29, 30. By numerically solving the corresponding equations for a consecutive nanoparticle radii, we obtained the radius-dependent resonance frequencies $\omega_{l}(R)$ and the associated damping rates $\Gamma_{l}(R)$ over an exceptionally wide range of particle sizes, from a few nanometers up to several hundred nanometers. Importantly, the size evolution of these quantities is not imposed through predefined scaling relations but emerges directly and self-consistently from the dispersion equation.
This methodology is applicable over an even broader size regime, extending in principle from the nanometer to the micrometer scale, provided that the employed dielectric response function accurately reproduces the metal’s optical properties in the spectral domain relevant to the multipolar LSP oscillation frequencies under consideration. The frequency window within which the dielectric function remains valid determines the admissible range of complex solutions $\omega_{l}(R)+i\Gamma_{l}(R)$. However, this constraint lies well outside the size range pertinent to the phenomena analyzed here and to typical experimental implementations.
2.2 Damping Rates of Multipolar LSP Oscillations vs Radius
@@
where $\Gamma_{l}^{r}(R)$ is the radiative damping rate (from the collisionless model), and $\Gamma^{nr}(R)\simeq\Gamma^{col}(R)=\gamma(R)/2$, if assumed to account for electron collisions, with:
@@
This effective collision rate appears explicitly in $\varepsilon_{in}$, incorporating both bulk ohmic losses $\gamma_{b}$ and the size-dependent contribution from surface scattering of electrons, where $v_{F}$ denotes the Fermi velocity and $C\approx 1$ for noble metals such as Au and Ag (e.g., 10, 30). The data from 10 (Figure 2) also phenomenologically incorporate the influence of interband transitions on the polarizability. However, the corresponding transition rate is not an explicit parameter in the model of the dielectric function.
Figure 2: (a) Total damping rates $\Gamma_{l}(R)$ (for $l$ = 1 - 6) calculated for gold NPs from DR (data from 28). (b) Close-up view showing that $\Gamma_{l}(R)\leq\gamma(R)/2$, suggesting non-additivity of radiative $\Gamma_{l}^{r}$ and nonradiative $\gamma(R)/2$ contributions.
@@
which may stem from a suppression of the nonradiative damping 58, 28. In particular, Sonnichsen et al. 58 attributed the suppression of nonradiative damping primarily to a reduction in interband transition processes.
@@
3 Foundations for Describing LSP Excitations as Quantum Plasmonic Quasi-Particles
The concept of modeling localized surface plasmon (LSP) excitations as quasi-particles in the excited state arises from their resonant behavior followed by relaxation, observable through the spectral properties of scattered or absorbed light. This approach establishes a bridge between the classical electromagnetic description based on Maxwell’s equations and a quantum-mechanical framework where plasmonic excitations are treated as quantized states of a collective many-electron system.
@@ 3 Foundations for Describing LSP Excitations as Quantum Plasmonic Quasi-Particles
Each multipolar mode $l$ corresponds to a discrete energy level of an excited quantum plasmonic quasi-particle (PQP), modeled as a set of uncoupled two-level subsystems $S_{l}$ undergoes transient decay to the ground state $|0\rangle$ after excitation.
@@
3.1 Closed System
In conventional CQED, models such as the Jaynes-Cummings Hamiltonian describe the interaction between a two-level atom and a single quantized mode of an external cavity field 24. This model assumes a closed system that conserves the number of quanta.
In the specific case of a closed confined subsystem $S_{l}$ of POP, it forms an integrated quantum system in which quantized TM polarized fields sourced in LSP oscillations and quasi-particle dynamics are intrinsically coupled (see Figure 3).
@@ 3.1 Closed System
The ground state $|0\rangle$ represents the stationary, non-oscillatory state of the POP, that is, the state in the absence of collective electron excitations. In this sense, the zero-energy state of the plasmonic quasi-particle is an internal ground state of the many-electron system, defined with respect to the collective charge-density oscillations supported by the nanoparticle geometry. The excited states $|l\rangle$ correspond to discrete collective oscillation modes of the electrons with eigenenergies $\hbar\omega_{l}(R)$.
In the formal bosonic quantization of plasmonic modes (e.g. 56), the full Hamiltonian includes a zero-point energy contribution of $\hbar\omega_{l}/2$. However, our description focuses on transitions between the excited states and the ground state, which involve energy differences of $\hbar\omega_{l}$. Consequently, the zero-point contribution cancels out and does not influence the decay dynamics or the self-quenching mechanism derived in this work. For this reason, the reduced Hamiltonian in Eq. (9), expressed in terms of the projection operators $|\outerproduct{l}{l}$, is entirely sufficient to describe the plasmonic quasi-particle energy structure relevant to emission and damping processes.
Such a self-consistent model provides a compact, physically grounded quantum framework for describing many-electron systems performing coherent collective motion. These systems can be viewed as micro- and macroscopic quantum structures (from tens to hundreds of nanometers in size), sharing characteristics related to the phenomenon of macroscopic quantum coherence demonstrated in superconducting circuits 12, 41 and recently awarded by the Nobel Prize.
3.2 Decay dynamics of Coherences and Populations in POP
In realistic conditions, the plasmonic quasi-particle system $S=\sum_{l}S_{l}$ interacts with its environment $E$, forming an open quantum system. Environmental coupling introduces dissipation and decoherence, leading to energy loss and phase randomization. These effects can be approximately described within the Lindblad master equation formalism under the Markov approximation 8.
@@ 3.2 Decay dynamics of Coherences and Populations in POP
The system evolves through two main decay channels: radiative decay, associated with photon emission, modeled by $L_{r,l}=\sqrt{2\Gamma_{l}^{reff}}\,\sigma_{l}^{-}$ and nonradiative decay, associated with internal losses (ohmic heating, electron–electron or electron–surface scattering), modeled by $L_{nr,l}=\sqrt{2\Gamma_{l}^{nr}}\,\sigma_{l}^{-}$. The radius dependence of the rates $\Gamma_{l}^{reff}(R)$ and $\Gamma_{l}^{nr}(R)$ with $\Gamma_{l}^{nr}=\Gamma^{nr}+\Delta_{l}$ follows from the classical modeling.
Solving the Lindblad equation for POP 31, 32 yields exponentially decaying populations and coherences:
@@
3.3 Spontaneous Emission and Electromagnetic Environment
@@
where $\omega_{21}$ is the transition frequency and $d_{12}$ is the dipole matrix element. This expression is sometimes analogously applied to describe the radiative damping of dipolar LSP modes (e.g. 44).
When the emitter is placed in an inhomogeneous medium, such as an optical cavity, the spontaneous emission rate is modified by the available EM states of the surrounding environment, altering the vacuum field mode structure and the local density of states. As a result, the emission rate can be enhanced or suppressed compared to that in free space (e.g., 40, 46. This fundamental sensitivity of spontaneous emission to the electromagnetic environment forms the basis of CQED (e.g., 39, 35, 11, 53, 48) and underpins emission control in structured photonic media.
Cavity quantum electrodynamics (CQED) and its plasmonic counterpart (PQED) investigate how optical and plasmonic cavities affect spontaneous emission via the Purcell effect, which quantifies enhancement (or suppression) of emission due to increased LDOS near the emitter’s resonance 19, 62, 21.
@@
For the conventional emitters in a plasmonic cavity, strong field confinement leads to very small $V_{eff}$, enabling high Purcell enhancement despite low $Q$. Plasmonic cavities of different geometries (e.g., bowtie, gap antenna) are essential tools in nanophotonics for squeezing light down to the nanoscale to unlock powerful light-matter interactions. However, in proximity to metallic surfaces, nonradiative decay processes become predominant, leading to quenching of the emission and a concomitant decrease in the lifetime of the emitter (e.g. 37, 2, 5, 18, 47, 51, 33, 21, 20)
4 Dual Functionality of Plasmonic Quasi-Particles